# Generic Antiholomorphic Polymonial Vector Fields

### v1

@@ 4. Analytic invariant
Let $\overline{P}{\partial\partial z}\in\xi_{k+1}$ and let $G$ be its $(k+2)$-nc tree. We define the $j$-th vertex $v_{j}$ to be at position $\exp\!\left({2i\pi(2j-1)2(k+2)}\right)$ on the circle. To lighten the notation, we will write $j$ instead of $v_{j}$.
@@
    We define a partial order on the edges. We say $\{j,m\}<\{j,\ell\}$ if $j<m<\ell$, $m<\ell<j$ or $\ell<j<m$. In other words, the edges attached to $j$ are ordered starting from $j$ and going around the circle in the positive direction.
@@
    Let $n_{j}$ be the incidence of the $j$-th vertex of $G$. Then there are $n_{j}-1$ sepal zones associated to this vertex (see Proposition 2.7). Let $e_{1}<\cdots<e_{n_{j}}$ be the edges attached to $j$ and $z_{e_{\ell}}$ be the singular point on $e_{\ell}$. Then we define
@@
    For $1\leq m\leq k$, there exist unique $21\leq j\leq k+$ 2 and $11\leq\ell\leq n_{j}-$ 1 such that $m=(n_{1}-1)+\cdots+(n_{j-1}-1)+\ell$, so that we define
    $a_{m}:={\rm segment}(j,\ell).$
    See Figure 7.
Figure 7. Line segments $a_{j}$ for the analytic invariants. Notice at the vertex 5, $a_{5}$ is on the right and $a_{6}$, on the left, because of the partial order on the edges.
The way we choose to order the sepal zones is not intrinsically important for our study. However, we must choose one to go forward.
@@
Let $P\in\Xi_{k+1}$, $G$ be the combinatorial invariant and $a_{j}$ be the line segments defined in Definition 4.1 using $G$. The analytic invariant of $\overline{P}{\partial\partial z}$ and ${1P}{\partial\partial z}$ is a vector $\eta\in\mathbb{H}^{k}$, where $\eta_{j}=\int_{a_{j}}P(z)\,{\rm d}z$.
@@
Let $k\in\mathbb{N}$ and $P_{1},P_{2}$ be two monic centred polynomials of degree $k+1$. If the two vector fields $\overline{P}_{1}{\partial\partial z}$ and $\overline{P}_{2}{\partial\partial z}$ are generic and have the same combinatorial and analytic invariants, then they are equal.
@@
We obtain a mapping $h\colon\mathbb{C}\to\mathbb{C}$ such that $h_{*}\bigl{(}{1P_{1}}{\partial\partial z}\bigr{)}={1P_{2}}{\partial\partial z}$ using the same construction as in the proof of Theorem 3.11, but now the mappings between the horizontal strips in the rectifying coordinates are replaced with the identity, since the analytic invariants are the same (i.e. the corresponding strips are equal). The rectifying coordinates and the identity are holomorphic and extend to the separatrices continuously, which defines $h$ as holomorphic bijection.
A conformal mapping of $\mathbb{C}$ on itself must be affine, and since the vector fields have the same combinatorial invariant, no rotation is allowed. Furthermore, both polynomials are centred, thus no translation is allowed. It follows that $h=id$ and $P_{1}=P_{2}$. ∎
Recall that a permutation $\rho(z)=\exp\!\left(2i\pi jk+2\right)$ acts on $(k+2)$-nc trees by a rotation of the vertices $\ell\mapsto\ell+j$ (mod $k+2$) and of the edges. Such a rotation will also change the order the analytic invariance by translating the indices $m\mapsto m+(n_{k+2-j+1}-1)+\cdots+({n_{k+2}}-1)$ (mod $k$).
@@
Suppose $\overline{P}_{j}{\partial\partial z}$ has invariants $(G_{j},\eta^{(j)})$ for $j=1,2$. If there exists a rotation $\rho$ that acts on $(G_{1},\eta^{(1)})$ so that $\rho.G_{1}=G_{2}$ and $\rho_{G_{1}}.\eta^{(1)}=\eta^{(2)}$, then we have $\rho_{*}\bigl{(}\overline{P}_{1}{\partial\partial z}\bigr{)}=\overline{P}_{2}{\partial\partial z}$, that is, with $\rho(z)=\lambda z$,
$\overline{P_{2}}(z)=\lambda\overline{\textstyle P_{1}\bigl{(}{z\lambda}\bigr{)}}.$
@@ 5. Realization
Let $k\in\mathbb{N}$. For every pair $(G,\eta)$, where $G$ is a $(k+2)$-noncrossing tree and $\eta\in\mathbb{H}^{k}$, there exists a unique monic centred polynomial $P$ such that ${1P}{\partial\partial z}$ has $(G,\eta)$ for its invariant. Moreover, $P$ is also the unique monic centred polynomial such that $\overline{P}{\partial\partial z}$ has $(G,\eta)$ for its invariant.
@@
We can quotient $\mathcal{A}_{k}\times\mathbb{H}^{k}$ by the relation $\sim$ defined by $(G_{1},\eta^{(1)})\sim(G_{2},\eta^{(2)})$ if and only if there exists a rotation $\rho$ of order $k+2$ such that $(G_{2},\eta^{(2)})=(\rho.G_{1},\rho_{G_{1}}.\eta^{(1)})$.
@@
Let $\Xi_{k+1}^{\prime}$ be the set of vector fields $\overline{P}{\partial\partial z}$ such that $P$ has degree $k+1$, is monic centred and normalized with $\arg(\overline{P(0)})\in\bigl{[}0,{2\pi k+2}\bigr{)}$. Then the moduli space of those generic vector fields is $\mathcal{A}_{k}\times\mathbb{H}^{k}/\sim$. It has $A^{r}(k+2)$ connected component, where $A^{r}(k+2)$ is given by Equation (7).
The proof of the theorem is similar to the realization construction found in [DES05] for a generic vector field $P{\partial\partial z}$. The combinatorial data and the analytic invariant allow us to construct a Riemann surface $\widetilde{M}$ with open sets of $\mathbb{C}$ in such a way that the vector field $1{\partial\partial z}$ is well-defined on $\widetilde{M}\setminus Z$, where $Z$ is a finite set. Then it is proved that $\widetilde{M}$ has the conformal type of the Riemann sphere and the push-forward of $1{\partial\partial z}$ will be ${1P}{\partial\partial z}$ for some polynomial $P$, after perhaps applying a Möbius transformation. To obtain $\overline{P}{\partial\partial z}$ is a bit more tricky, since the push-forward of an antiholomorphic vector field is not, in general, antiholomorphic. The trick is to push-forward a well-chosen log-harmonic function: if $h\colon\widetilde{M}\to\hat{\mathbb{C}}$ is the appropriate isomorphism, then we will have $h_{*}\bigl{(}|h^{\prime}|^{-2}{\partial\partial\zeta}\bigr{)}=\overline{P}{\partial\partial z}$.
@@ 5.1. Construction of the Riemann surface
Consider a $(k+2)$-nc tree $G$ and a vector $\eta\in\mathbb{H}^{k}$. Recall that the dual $G^{*}$ is also a $(k+2)$-nc tree. If we embed both $G$ and $G^{*}$ in the same closed disk $\overline{\mathbb{D}}$ so that the $j$-th vertex of $G$ is mapped to $e_{2j-1}$ and the $j$-th vertex of $G^{*}$, on $e_{2j-2}$, where $e_{\ell}:=\exp\!\left(2i\pi{\ell 2(k+2)}\right)$, they divide the open disk in $43k+$ 4 connected and simply connected components, provided the embedding is chosen so that two edges intersect at most once. We will call the intersection of two edges a singular point, which divides those edges in fours arcs that we will call the separatrices. By Proposition 2.7, there are $42k+$ 4 components adjacent to only one singular point, which we call the petal zones, and $k$ components adjacent to two singular points, the sepal zones.
@@
The sepal zones are ordered by $G$ from 1 to $k$. The $j$-th sepal zone can be mapped by a homeomorphism to a horizontal strip of width $\Im\eta_{j}$, which we can choose in a way that it extends to the boundary so as to map the singular points, say $z_{0}$ and $z_{1}$, on 00 and $\eta_{j}$. There are two orientations for this; to choose the correct one, note that a sepal zone has an end on a vertex $e_{\ell^{-}}$ of $G$ and the other end on a vertex $e_{\ell^{+}}$ of $G^{*}$. We choose a homeomorphism that maps a horizontal line in the strip oriented from left to right to a curve in $\mathbb{D}$ from $e_{\ell^{-}}$ to $e_{\ell^{+}}$. Then the four separatrices in $\mathbb{D}$ are forced to be mapped on horizontal lines on each side of 00 and $\eta_{j}$ in the strip.
@@
Let $Z_{0}$ be a singular point and $S_{1},S_{2},S_{3},S_{4}$ the four separatrices attached to it, with $S_{1}$ an outgoing separatrix (any one of the two) and $S_{2},S_{3},S_{4}$ going anti-clockwise around $Z_{0}$. Then $Z_{0}$ appears on the upper boundary of two charts, and lower boundary of two charts. For $m\in\{1,2,3,4\}$, we define $D_{m}$ as a upper (for $m$ odd) or lower (for $m$ even) half-disk from $S_{m}$ to $S_{m+1}$ (with $S_{5}=S_{1}$) with small enough radius centred at $Z_{0}$ and contained in the chart. If we glue the $D_{m}$’s along the separatrices, we obtain a neighbourhood $D_{Z_{0}}$ of $Z_{0}$. Let
@@
This defines an isomorphism $\psi_{Z_{0}}\colon D_{Z_{0}}\setminus\{Z_{0}\}\to D(0,r)\setminus\{0\}$ for some $r>0$. We can extend $\psi_{Z_{0}}$ to $Z_{0}$ and obtain a chart $(D_{Z_{0}},\psi_{Z_{0}})$ of $Z_{0}$. See Figure 8. We can verify that $(\psi_{Z_{0}})_{*}(1{\partial\partial Z})(z)={12z}{\partial\partial z}$, so that the dynamics is that of a simple saddle node.
@@
We will label vertices of $G^{*}$ by $1^{+},\ldots,(k+2)^{+}$, with the $j$-th vertex positioned at $\exp\!\left({2i\pi jk+2}\right)$. Let $U$ be a neighbourhood of infinity. For a strip $B$, $B\cap U$ has two components that we call ends. Let $\mathcal{E}_{j}^{+}$ (resp. $\mathcal{E}_{j}^{-}$) the end that lies at the $j$-th (resp. $j^{+}$-th) vertex. See Figure 9(a).
Starting with the petal zone $H_{1}^{+}$ between the vertices $1^{+}$ and $1$, we glue the positive ends at $1^{+}$ (if any) to $H_{1}^{+}\cap U$ and we translate by the the analytic invariant of each of those ends. We obtain an open set $U_{1}^{+}$ of $\mathbb{C}$. We define $\psi_{1}^{+}\colon U_{1}^{+}\to\mathbb{C}$ by $\psi_{1}^{+}(Z)=Z^{-{1k+2}}$, with $\arg(Z)\in\bigl{(}-{\pi 2},{3\pi 2}\bigr{)}$.
For the next petal zone $H_{1}^{-}$ (between $1$ and $2^{+}$), we glue negative ends at $1$ (if any) to $H_{1}^{-}\cap U$ and we translate by the analytic invariant of the positive ends at $1^{+}$ and by minus the analytic invariant of negative ends at 1 to obtain an open set $U_{1}^{-}$. We then define $\psi_{1}^{-}\colon U_{1}^{-}\to\mathbb{C}$ by $\psi_{1}^{-}(Z)=Z^{-{1k+2}}$ with $\arg(Z)\in\bigl{(}{\pi 2},{5\pi 2}\bigr{)}$.
We define $U_{j}^{\pm}$ and $\psi_{j,\infty}^{\pm}$ for every petal zones. See Figure 9(b). This defines a conformal map $\psi_{\infty}$ on $U$ that sends $U$ on some punctured neighboorhood of 00. We can then add $\infty$ to $M$ and $\psi_{\infty}$ extends as a holomorphic chart of $\infty$ on $M$.
(a) Neighbourhood of infinity represented in $\overline{\Delta}$, given a noncrossing tree
@@
Let $\widetilde{M}$ be the Riemann surface obtained from $M$ after adding the singular points $\{Z_{j}\}_{1\leq j\leq k+1}$ and $\infty$. It is a compact manifold with no boundary. We can compute the Euler characteristic with the following triangulation: the vertices are the $k+1$ singular points and $\infty$, the edges are the $4(k+1)$ separatrices, and the faces are the $2(k+2)$ half-planes and the $k$ strips. We have $k+2$ vertices, $4(k+1)$ edges and $43k+$ 4 faces, so that $\chi_{\widetilde{M}}=2$. Therefore, $\widetilde{M}$ is a topological sphere and thus, by Riemann’s Uniformization Theorem, it has the conformal type of the Riemann’s sphere ${}^\hat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$.

### v2

@@ 4. Analytic invariant
Let $\overline{P}{\partial\partial z}\in\Xi_{k+1}^{\prime}$ and let $G$ be its $(k+2)$-nc tree. Recall that the vertex $j$ (mod $k+2$) is at position $\exp\!\left({2i\pi(2j-1)2(k+2)}\right)$ on the circle.
@@
    We define a partial order on the edges. We say $\{j,m\}<\{j,\ell\}$ if $j<m<\ell$, with $m,l\in\{j+1,j+2,...,j+k+1\}$. In other words, the edges attached to $j$ are ordered starting from $j$ and going around the circle anti-clockwise.
@@
    We define an order on the sepal zones. Let $n_{j}$ be the incidence of the $j$-th vertex of $G$. Then there are $n_{j}-1$ sepal zones associated to this vertex (see Proposition 2.10). For $1\leq m\leq k$, there exist unique $21\leq j\leq k+$ 2 and $11\leq\ell\leq n_{j}-$ 1 such that $m=(n_{1}-1)+\cdots+(n_{j-1}-1)+\ell$. This means that the $m$-th sepal zone is the $\ell$-th sepal zone of the $j$-th vertex. See Figure 7.
  3. (3)
    Let $e_{1}<\cdots<e_{n_{j}}$ be the edges attached to $j$ and $z_{e_{\ell}}$ be the singular point on $e_{\ell}$. Then we define
@@
    For the $m$-th sepal zone, we define the line segment $a_{m}:=\text{segment}(j,l)$, where $j$ and $\ell$ are as in (2).
Figure 7. Line segments $a_{j}$ for the analytic invariants. Notice that at the vertex 5, $a_{5}$ is on the right and $a_{6}$, on the left, because of the partial order on the edges.
The way we choose to order the sepal zones is not intrinsically important for our study, however, we must choose one to go forward.
@@
Let $\overline{P}\in\Xi_{k+1}^{\prime}$, $G$ be the combinatorial invariant and $a_{j}$ be the line segments defined in Definition 4.1 using $G$. The analytic invariant of $\overline{P}{\partial\partial z}$ and ${1P}{\partial\partial z}$ is the vector $\eta\in\mathbb{H}^{k}$, with the $j$-th component $\eta_{j}=\int_{a_{j}}P(z)\,{\rm d}z$.
@@
Let $k\in\mathbb{N}$ and $P_{1},P_{2}$ be two monic centred polynomials of degree $k+1$. If the two vector fields $\overline{P}_{1}{\partial\partial z}$ and $\overline{P}_{2}{\partial\partial z}$ are generic and have the same combinatorial and analytic invariants, then they are analytically equivalent.
@@
We obtain a mapping $h\colon\mathbb{C}\to\mathbb{C}$ such that $h_{*}\bigl{(}{1P_{1}}{\partial\partial z}\bigr{)}={1P_{2}}{\partial\partial z}$ using the same construction as in the proof of Theorem 3.10, but now the mappings between the horizontal strips in the rectifying coordinates are replaced with the identity, since the analytic invariants are the same (i.e. the corresponding strips are equal). The rectifying coordinates and the identity are holomorphic and extend to the separatrices continuously, which defines $h$ as holomorphic bijection. ∎
Recall that a rotation $\rho(z)=\exp\!\left(2i\pi jk+2\right)$ acts on $(k+2)$-nc trees by a rotation of the vertices $\ell\mapsto\ell+j$ (mod $k+2$) and of the edges. Such a rotation will also change the order of the components of the analytic invariant by translating the indices $m\mapsto m+(n_{k+2-j+1}-1)+\cdots+({n_{k+2}}-1)$ (mod $k$).
@@
Suppose a generic $\overline{P_{j}}{\partial\partial z}$ has invariants $(G_{j},\eta^{(j)})$ for $j=1,2$. If there exists a rotation $\rho$ that acts on $(G_{1},\eta^{(1)})$ so that $\rho.G_{1}=G_{2}$ and $\rho_{G_{1}}.\eta^{(1)}=\eta^{(2)}$, then $\rho_{*}\bigl{(}\overline{P_{1}}{\partial\partial z}\bigr{)}$ and $\overline{P_{2}}{\partial\partial z}$ are analytically equivalent, where, with $\rho(z)=\lambda z$,
$\rho_{*}\bigl{(}\overline{P}_{1}{\partial\partial z}\bigr{)}=\lambda\overline{\textstyle P_{1}\bigl{(}{z\lambda}\bigr{)}}.$
@@ 5. Realization
Let $k\in\mathbb{N}$. For every pair $(G,\eta)$, where $G$ is a $(k+2)$-noncrossing tree and $\eta\in\mathbb{H}^{k}$, there exists a monic centred generic polynomial $P$ of degree $k+1$ such that $\overline{P}{\partial\partial z}$ has $(G,\eta)$ for its invariant.
@@
If we allow homeomorphisms to rotate points at infinity, this corresponds to orientation-preserving topological orbital equivalence. In this case, we can quotient $\mathcal{A}_{k}\times\mathbb{H}^{k}$ by the relation $\sim$ defined by $(G_{1},\eta^{(1)})\sim(G_{2},\eta^{(2)})$ if and only if there exists a rotation $\rho$ of order $k+2$ such that $(G_{2},\eta^{(2)})=(\rho.G_{1},\rho_{G_{1}}.\eta^{(1)})$.
@@
Let $\widetilde{\Xi}_{k+1}^{\prime}$ be the set of generic vector fields $\overline{P}{\partial\partial z}$ such that $P$ has degree $k+1$, is monic centred and normalized with $\arg(\overline{P(0)})\in\bigl{[}0,{2\pi k+2}\bigr{)}$. Then the moduli space of those generic vector fields is $\mathcal{A}_{k}\times\mathbb{H}^{k}/{\sim}$. It has $A^{r}(k+2)$ connected component, where $A^{r}(k+2)$ is given by Equation (8).
The proof of the theorem is similar to the realization construction found in [DES05] for a generic vector field $P{\partial\partial z}$. The combinatorial data and the analytic invariant allow us to construct a Riemann surface $\widetilde{M}$ with open sets of $\mathbb{C}$ in such a way that the vector field $1{\partial\partial z}$ is well-defined on $\widetilde{M}\setminus Z$, where $Z$ is a finite set. Then it is proved that $\widetilde{M}$ has the conformal type of the Riemann sphere and the push-forward of $1{\partial\partial z}$ will be ${1P}{\partial\partial z}$ for some polynomial $P$, after perhaps applying a Möbius transformation. To obtain $\overline{P}{\partial\partial z}$ is a bit more tricky, since the push-forward of an antiholomorphic vector field is not, in general, antiholomorphic. The trick is to push-forward a well-chosen log-harmonic function: if $h\colon\widetilde{M}\to\hat{\mathbb{C}}$ is the isomorphism with $h_{*}\bigl{(}1{\partial\partial\zeta}\bigr{)}={1P}{\partial\partial z}$, then we will have $h_{*}\bigl{(}|h^{\prime}|^{-2}{\partial\partial\zeta}\bigr{)}=\overline{P}{\partial\partial z}$.
@@ 5.1. Construction of the Riemann surface
Consider a $(k+2)$-nc tree $G$ and a vector $\eta\in\mathbb{H}^{k}$. Recall that the dual $G^{*}$ is also a $(k+2)$-nc tree. If we embed both $G$ and $G^{*}$ in the same closed disk $\overline{\mathbb{D}}$ so that the $j$-th vertex of $G$ is mapped to $e_{2j-1}$ and the $j$-th vertex of $G^{*}$, on $e_{2j-2}$, where $e_{\ell}:=\exp\!\left(2i\pi{\ell 2(k+2)}\right)$, they divide the open disk in $43k+$ 4 connected and simply connected components (see Proposition 2.10), provided the embedding is chosen so that two edges intersect at most once. We will call the intersection of two edges a singular point, which divides those edges in four arcs that we will call the separatrices. By Proposition 2.10, there are $42k+$ 4 components adjacent to only one singular point, which we call the petal zones, and $k$ components adjacent to two singular points, the sepal zones.
@@
The sepal zones are ordered by $G$ from 1 to $k$ according to Definition 4.1. The $j$-th sepal zone can be mapped by a homeomorphism to a horizontal strip $B_{j}$ of width $\Im\eta_{j}$, which we can choose in a way that it extends to the boundary so as to map the singular points, say $z_{0}$ and $z_{1}$, on 00 and $\eta_{j}$. There are two orientations for this; to choose the correct one, note that a sepal zone has an end on a vertex $e_{\ell^{-}}$ of $G$ and the other end on a vertex $e_{\ell^{+}}$ of $G^{*}$. We choose a homeomorphism that maps a horizontal line in the strip oriented from left to right to a curve in $\mathbb{D}$ from $e_{\ell^{-}}$ to $e_{\ell^{+}}$. Then the four separatrices in $\mathbb{D}$ are forced to be mapped on horizontal lines on each side of 00 and $\eta_{j}$ in the strip.
@@
Let $Z_{0}$ be a singular point and $S_{1},S_{2},S_{3},S_{4}$ the four separatrices attached to it, with $S_{1}$ an outgoing separatrix (any one of the two) and $S_{2},S_{3},S_{4}$ going anti-clockwise around $Z_{0}$. Then $Z_{0}$ appears on the upper boundary of two charts, and lower boundary of two charts. For $m\in\{1,2,3,4\}$, we define $D_{m}$ as a upper (for $m$ odd) or lower (for $m$ even) half-disk from $S_{m}$ to $S_{m+1}$ (with $S_{5}:=S_{1}$) with small enough radius centred at $Z_{0}$ and contained in the chart. If we glue the $D_{m}$’s along the separatrices, we obtain a neighbourhood $D_{Z_{0}}$ of $Z_{0}$. Let
@@
This defines an isomorphism $\psi_{Z_{0}}\colon D_{Z_{0}}\setminus\{Z_{0}\}\to D(0,r)\setminus\{0\}$ for some $r>0$. By Riemann’s Removable Singularity Theorem, we can extend $\psi_{Z_{0}}$ to $Z_{0}$ and obtain a chart $(D_{Z_{0}},\psi_{Z_{0}})$ of $Z_{0}$. See Figure 8. We can verify that $(\psi_{Z_{0}})_{*}(1{\partial\partial Z})(z)={12z}{\partial\partial z}$, so that the dynamics is that of a simple saddle node.
@@
We will label vertices of $G^{*}$ by $1^{+},\ldots,(k+2)^{+}$, with the $j$-th vertex positioned at $\exp\!\left({2i\pi(j-1)k+2}\right)$. Let $U$ be a neighbourhood of infinity. For a strip $B$, $B\cap U$ has two components that we call ends. (See Figure 9(a).) Let $E_{j}^{\pm}=\{1\leq m\leq k\mid B_{m}\text{ has an end at vertex }j^{\pm}\}$ (it may be empty). For each $m\in E_{j}^{\pm}$, we define $\mathcal{E}_{m}^{\pm}$ as the end of $B_{m}$ at vertex $j^{\pm}$, where the strips $B_{1},\ldots,B_{k}$ are ordered following the sepal zones.
Starting with the petal zone $H_{1}^{+}$ between the vertices $1^{+}$ and $1$, we glue the positive $\mathcal{E}_{m}^{+}$ ($m\in E_{1}^{+}$) ends at $1^{+}$ (if any) to $H_{1}^{+}\cap U$ and we translate by $\xi_{1}^{+}=\sum_{m\in E_{1}^{+}}\eta_{m}$, see Figure 9(b). We obtain an open set $U_{1}^{+}$ of $\mathbb{C}$. We define $\psi_{1}^{+}\colon U_{1}^{+}\to\mathbb{C}$ by $\psi_{1}^{+}(Z)=Z^{-{1k+2}}$, with $\arg(Z)\in\bigl{(}-{\pi 2},{3\pi 2}\bigr{)}$.
For the next petal zone $H_{1}^{-}$ (between $1$ and $2^{+}$), we glue negative ends $\mathcal{E}_{m}^{-}$ ($m\in E_{1}^{-}$) at $1$ (if any) to $H_{1}^{-}\cap U$ and we translate by $\xi_{1}^{-}=\sum_{m\in E_{1}^{+}}\eta_{m}-\sum_{m\in E_{1}^{-}}\eta_{m}$ to obtain an open set $U_{1}^{-}$. We then define $\psi_{1}^{-}\colon U_{1}^{-}\to\mathbb{C}$ by $\psi_{1}^{-}(Z)=Z^{-{1k+2}}$ with $\arg(Z)\in\bigl{(}{\pi 2},{5\pi 2}\bigr{)}$. The translation by $\xi_{1}^{-}$ is necessary so that when $U_{1}^{+}$ and $U_{1}^{-}$ are glued, the $\psi_{1}^{\pm}$ extend to an injective map on the quotient into a connected open set of $\mathbb{C}$.
For the general case, we glue the positive (resp. negative) ends at $j$ (resp. $j^{+}$) to $H_{j}^{+}\cap U$ (resp. $H_{j}^{-}\cap U$) and we translate by
$\xi_{j}^{+}=\sum_{\ell=1}^{j-1}\Bigg{[}\sum_{m\in E_{\ell}^{+}}\eta_{m}-\sum_{m\in E_{\ell}^{-}}\eta_{m}\Bigg{]}+\sum_{m\in E_{j}^{+}}\eta_{m}\quad\left(\hbox{resp.~{}}\xi_{j}^{-}=\sum_{\ell=1}^{j}\Bigg{[}\sum_{m\in E_{\ell}^{+}}\eta_{m}-\sum_{m\in E_{\ell}^{-}}\eta_{m}\Bigg{]}\right).$
We obtain an open set $U_{j}^{+}$ (resp. $U_{j}^{-}$) of $\mathbb{C}$. We define $\psi_{j,\infty}^{+}\colon U_{j}^{+}\to\mathbb{C}$ (resp. $\psi_{j,\infty}^{-}\colon U_{j}^{-}\to\mathbb{C}$) by $\psi_{j}^{\pm}(Z)=Z^{-{1k+2}}$ with $\arg(Z)\in\bigl{(}(4j-5){\pi 2},(4j-1){\pi 2}\bigr{)}$ (resp. in $\bigl{(}(4j-3){\pi 2},(4j+1){\pi 2}\bigr{)}$). See Figure 9(b). This defines a conformal map $\psi_{\infty}$ on $U$ that sends $U$ on some punctured neighboorhood of 00. We can then add $\infty$ to $M$ and $\psi_{\infty}$ extends as a holomorphic chart of $\infty$ on $M$.
(a) Neighbourhood of infinity represented in $\overline{\Delta}$, given a noncrossing tree, with $\mathcal{E}_{j}^{\pm}$ representing the ends of sepal zones
@@
Let $\widetilde{M}$ be the Riemann surface obtained from $M$ after adding the singular points $\{Z_{j}\}_{1\leq j\leq k+1}$ and $\infty$. It is a compact manifold with no boundary. We can compute the Euler characteristic with the following triangulation: the vertices are the $k+1$ singular points and $\infty$, the edges are the $4(k+1)$ separatrices, and the faces are the $2(k+2)$ half-planes and the $k$ strips. We have $k+2$ vertices, $4(k+1)$ edges and $43k+$ 4 faces, so that $\chi_{\widetilde{M}}=2$. Therefore, $\widetilde{M}$ is a topological sphere and thus, by Riemann’s Uniformization Theorem, it has the conformal type of the Riemann sphere ${}^\hat{\mathbb{C}}=\mathbb{C}\cup\{\infty\}$.